# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

### v2

@@ III-A Preliminary
In the initialization stage, typical algorithms such as HWF, SPARTA and TP usually aim at producing a good estimate $\mathbf{z}$ falling into the $\delta$ neighborhood of the target signal, i.e.,
@@
which is the prerequisite of the refinement stage [19, 20, 29]. In the analysis of gESP, we follow the same line and focus on obtaining the estimate satisfying (32).
@@
Let $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ be the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. the structure function $s(p)$ is given by
@@
The structure function, $s(p)$, quantifies the structural information of a signal $\mathbf{x}$ by measuring the energy distribution across its components. Specifically, $s(p)$ represents the proportion of the total signal energy captured by the $p$ largest magnitude components. Therefore, the behavior of $s(p)$ as $p$ varies directly reflects how energy is concentrated or dispersed within the signal, thus characterizing its structure.
@@
    $s(p)$ is monotonously decreasing with respect to $p$. Furthermore,
@@
    which implies $ps(p)$ is monotonously increasing with respect to $p$. Furthermore,
@@ III-B Theoretical results for sample complexity of gESP
For any numerical constant $0<\delta<1$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ with probability at least $1-4e^{-c_{\delta}m}$. The corresponding required sample complexity depends on an input parameter $p\in[k]$; for a general $p$, it is given by
@@
The state-of-the-art result [28] is
@@
Since $s(p)\geq 1$ (see (34)), (41) implies $s(p)=\Theta(1)$ for some $p\in[k]$. Then, (42) requires $p\leq c\lceil\sqrt{k}\rceil$ for some constant $c$. Moreover, if $s(p)=\Theta(1)$ holds for some fixed $p\leq c\lceil\sqrt{k}\rceil$, then $s(\lceil\sqrt{k}\rceil)=\Theta(1)$ since $s(p)$ is monotonously decreasing and larger than $1$. Therefore, to achieve the information-theoretical bound, gESP requires $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
This result provides the general sample complexity for gESP. However, in the case where $s(p)$ is unknown, determining the optimal $p$ is challenging. One compromise is to execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn. Then we shall obtain a set containing $k$ estimates of the target signal $\mathbf{x}$. It can be guaranteed that at least one estimate falls into the $\delta$-neighborhood of $\mathbf{x}$ with the same sample complexity requirement as (39). The formal statement is in the following corollary.
@@
Execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn, and collect all the estimates into a set $\mathcal{X}$. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, there exists at least one estimate $\mathbf{z}\in\mathcal{X}$ that satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ when
@@
When multiple executions of the algorithm with varying $p$ are impractical due to computational constraints, we are compelled to explore the sample complexity for specific, analytically tractable choices of $p$. A natural starting point for such an analysis is the intuitive choice of setting $p=k$. While this can be viewed as a special case of Theorem 1, we demonstrate that a more refined analysis allows us to derive a sharper result, as presented in the following corollary.
@@
Take $p=k$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number of samples $m$ satisfies
@@
Take $p=\lceil\sqrt{k}\rceil$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number $m$ of samples satisfies
@@
One one hand, the result (45) cannot be worse than (44) since
@@
Moreover, there exist some signals that make the less-than sign strictly holds. Therefore, (45) is indeed tighter than (44). See Section IV-B for an in-depth discussion.
On the other hand, compared to the state of the art, the results (44) and (45) are superior for certain signals. In other words, even though $s(p)$ is unknown as well as executing Algorithm 1 for multiple times is infeasible, our algorithm taking $p=k$ or $\lceil\sqrt{k}\rceil$ can perform better compared with existing algorithms. Discussion IV-A gives an in-depth discussion.
@@ IV-A Broader Conditions for Optimal sample complexity
As has been discussed in Remark 1 and Remark 2, gESP makes a significant improvement over the state-of-the-art result, if the structural function $s(p)$ is known. This subsection further explores this superiority, demonstrating that gESP achieves the optimal sample complexity for a broader class of signals.
@@
requires $s(1)=\Theta(1)$, which is a more strict condition. The improvement results from taking $p$ entries in Step 1 into consideration, thus relaxing the condition on $x_{(1)}$ to $x_{(1)},x_{(1)},\dots,x_{(p)}$. Fig. 1 gives a comprehensive illustration.
@@
Moreover, even in the case where $s(p)$ is unknown, our results (Corollary 2 and 3) can also perform better in many cases. In the following, we set $p=k$ and consider Corollary 2 for a simple discussion, although we can prove the strict superiority of the result derived in Corollary 3 (see Section IV-B). The result in Corollary 2 is
@@
Therefore, (48) outperforms the state-of-the-art result (47) once $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$. Here, we offer an example to demonstrate the existence of such signals.
@@
Assume that both $\sqrt{k}$ and $\sqrt[6]{k}$ are integers [^4] In fact, we can also construction such examples even though this assumption does not hold. However, omitting it will lead to unnecessary complications in the construction (e.g., involving terms like $\lceil\sqrt{k}\rceil\pm 1$). The same phenomenon also occurs in Example 2. For the sake of simplicity, we make this assumption. The target signal $\mathbf{x}$ satisfies
$|x_{(i)}|^{2}=\begin{cases}\frac{\|\mathbf{x}\|^{2}}{\sqrt{k}},&i=1,\\ \frac{\|\mathbf{x}\|^{2}}{\sqrt{k}-1}\times\left(\frac{1}{\sqrt[6]{k}}-\frac{1}{\sqrt{k}}\right),&i=2,\cdots,\sqrt{k},\\ \frac{\|\mathbf{x}\|^{2}}{k-\sqrt{k}}\times\left(1-\frac{1}{\sqrt[6]{k}}\right),&i=\sqrt{k}+1,\cdots,k.\end{cases}$
In this case, $s(1)=\sqrt{k}$ and $s(\sqrt{k})=\sqrt[6]{k}$, and one can thus find (48) outperforms (47).
Interestingly, when the condition $s(\lceil{\sqrt{k}}\rceil)=\Theta(1)$ holds, our result (48) also achieves the information-theoretic bound (2). Significantly, even without prior knowledge of the signal structure $s(p)$, this optimality is attainable for the same class of signals in the case $s(p)$ is known, simply by setting the parameter $p=k$.
@@ IV-B Strict superiority of Corollary 3 over Corollary 2
is better than (48) in Section III-B. In this section, we provide one certain example to illustrate this strict superiority.
@@
Assume that both $\sqrt{k}$ and $\sqrt[4]{k}$ are integers. Construct the target signal $\mathbf{x}$ as
@@
This demonstrate the improvement in Corollary 3.
@@ V Proof
The proofs of Theorem 1 and Corollary 13 are mainly based on the analysis for each step of gESP. In the following, we provide several propositions with respect to each step respectively.
@@
    When the number of samples satisfies $m\geq Cp^{2}s^{2}(p)\log n$, it holds with probability exceeding $1-2e^{-cm}$ that
@@
  • Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
@@
    Suppose the set $S^{0}$ contains $p^{\prime}$ indices and satisfies (49), then for any constant $0<\mu_{1}<1$, the inequality
    $|\mathbf{x}^{*}\mathbf{e}^{0}|\geq\mu_{1}\|\mathbf{x}_{S^{0}}\|$ (50)
    holds with probability at least $1-2e^{-c_{\mu_{1}}m}$ with the proviso that $m\geq{C_{\mu_{1}}}{p^{\prime}s^{2}(p)}\log n$, where $c_{\mu_{1}}$ and $C_{\mu_{1}}$ are constants related to $\mu_{1}$.
@@
    It may seem unusual to introduce a new parameter $p^{\prime}$ in this proposition, given that $S^{0}$ contains $p$ indices in Algorithm 1. However, the proposition holds even when $p^{\prime}=p$, making it a generalized version. We emphasize that this generalization will be useful in our proofs of Corollary 2 and 3, where we do not have access to the optimal choice of $p$.
  • Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in modulus.
@@
    where $\gamma\in(0,1)$ is a constant. Moreover, suppose that (49) holds and $\mathbf{e}^{0}$ satisfies (50). Then, if the number of samples satisfies $m\geq C_{\mu_{1}}\gamma^{-2}{ks(p)}\log n$, it holds with probability at least $1-4e^{-c_{\mu_{1},\gamma}m}$ that
@@
    where $\mu_{1}$ is the hyperparameter given in (50) and $c_{\mu_{1},\gamma}$ is the constant depending on $\mu_{1}$ and $\gamma$. In other words, gESP shall select all the indices of $S_{\gamma}$ in Step 3.
  • Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
@@
    Suppose $S^{1}\supseteq S_{\gamma}$. For any constant $0<\delta<1$ satisfying
    $\delta^{2}\geq 2-2\sqrt{\left(1-\frac{\gamma^{2}}{4}\right)^{3}}+\frac{\gamma^{2}}{4},$ (51)
    Step 4 produces a signal estimation $\mathbf{z}$ falling into the $\delta$-neighborhood of $\mathbf{x}$, i.e.,
@@
    with probability exceeding $1-2e^{-c_{\delta}m}$ providing that $m\geq{C_{\delta}}{k}\log n$, where $c_{\delta}$ and $C_{\delta}$ are numerical constants depending on $\delta$.
Now we are proceeding to prove the theorems in Section III.
@@
Collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the sample complexity is
@@
and the overall probability is $1-e^{-cm}$. Since we take the optimal value for $p$ as
@@
For each fixed $p\in[k]$, collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$ when
@@
To ensure that at least one estimate $\mathbf{z}$ satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we can apply the union bound for all $p\in[k]$. Then the sample complexity can be reduced to
@@
and the probability is at least $1-ke^{-cm}$. It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$. Thus the proof is complete. ∎
@@
(c) Exponential decaying signal. $k$ = 10
Figure 3: Performance comparison of relative error and fraction of recovered support as a function of sampling ratio.
@@
where $(a)$ is because $k\geq p$, and $(b)$ holds via applying Proposition 1 on the set $S^{\prime}$ when $\geq Cp^{2}s^{2}(p)\log n$.
@@
the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$. This result also holds for all $p\in[k]$ since we can take the union bound, and the probability will be $1-ke^{-cm}$. It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$.
@@
The remaining thing is to show that (56) is equivalent to (44).
As stated following Definition 2, $ps(p)$ is monotonously increasing, and $s(p)$ is monotonously decreasing. Both of them are positive. Accordingly, we can make an illustrative plot for the two functions in (56); see Fig. 2.
@@
From Lemma 1, we know that both the cases imply the same sample complexity
@@
The proof is similar to the proof of Corollary 2. Assume any fixed $p\in[\lceil{\sqrt{k}}\rceil]$. The sample complexity for Step 1 is still $\Omega(p^{2}s^{2}(p)\log n)$, and applying Proposition 24 on the rest steps with $p^{\prime}=\lceil{\sqrt{k}}\rceil$ finally leads to

### v3

@@ III-A Preliminary
In the initialization stage, typical algorithms, such as HWF, SPARTA and TPM, usually aim to produce a good estimate $\mathbf{z}$ falling into the $\delta$-neighborhood of the target signal, i.e.,
@@
which is the prerequisite of the refinement stage [4, 27, 34]. In the analysis of gESP, we follow the same line and focus on obtaining an estimate satisfying (32).
@@
Let $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ be the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes. The structure function $s(p)$ is given by
@@
The structure function, $s(p)$, quantifies the structural information of a signal $\mathbf{x}$ by measuring the energy distribution across its components. Specifically, $s(p)$ represents the inverse of the proportion of the total signal energy captured by the $p$ largest magnitude components. Therefore, the behavior of $s(p)$ as $p$ varies directly reflects how energy is concentrated or dispersed within the signal, thus characterizing its structure.
@@
    $s(p)$ is monotonically decreasing with respect to $p$. Furthermore,
@@
    which implies $ps(p)$ is monotonically increasing with respect to $p$. Furthermore,
@@
Finally, to facilitate rigorous theoretical analysis and decouple statistical dependencies among different iterations, we employ a standard sample splitting strategy [13, 5] in our proof. Specifically, the measurement set $\mathcal{D}=\{(a_{i},y_{i})\}_{i=1}^{m}$ is randomly partitioned into $4$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{4}$, where each subset is used for a single step of our algorithm. Here we emphasize that sample splitting is purely a theoretical artifact used to simplify the probabilistic analysis. In our practical implementation (Algorithm 1) and numerical experiments, we utilize the full dataset $\mathcal{D}$ at every step to maximize the sample efficiency. Since the number of steps is small and fixed (i.e., $4$), splitting the data only affects the sample complexity by a constant factor, leaving the order of the convergence rate unchanged.
@@ III-B Theoretical results for sample complexity of gESP
For any constant $0<\delta<1$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ with probability at least $1-n^{-c}$. The corresponding required sample complexity depends on an input parameter $p\in[k]$; for a general $p$, it is given by
@@
The state-of-the-art result [33] is the special case $p=1$ of our result, which is
@@
The statistical-to-computational gap persists for signals with maximally delocalized energy. Specifically, the worst-case sample complexity of $\Omega(k^{2}\log n)$ is required if and only if the signal is almost “flat”, i.e., $s(p)=\Theta\left(\frac{k}{p}\right)$ for all $p\in[1,k]$.
Remark 3.
@@
Since $s(p)\geq 1$ (see (34)), (41) implies $s(p)=\Theta(1)$ for some $p\in[k]$. Then, (42) requires $p\leq\lceil c\sqrt{k}\rceil$ for some constant $c$. Moreover, if $s(p)=\Theta(1)$ holds for some fixed $p\leq\lceil c\sqrt{k}\rceil$, then $s(\lceil\sqrt{k}\rceil)=\Theta(1)$ since $s(p)$ is monotonically decreasing and larger than $1$. Therefore, to achieve the information-theoretical bound (2), gESP requires $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
This result provides the general sample complexity for gESP. Interestingly, this result depends solely on the cumulative energy profile $s(p)$, imposing no lower bound requirement on the minimum non-zero magnitude $|x_{(k)}|$. This is due to that the assumption (11) is required for exact support recovery, whereas our analysis only targets an $\ell_{2}$-distance bound ($\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$), which, therefore, can be met under a relaxed requirement. A more specific explanation is provided below Proposition 3.
In the case where $s(p)$ is unknown, determining the optimal $p$ is challenging. One compromise is to execute Algorithm 1 $k$ times with $p$ taking each value in $[k]$. This yields a set containing $k$ estimates of the target signal $\mathbf{x}$. Naturally, at least one estimate falls into the $\delta$-neighborhood of $\mathbf{x}$ with the same sample complexity requirement as (39), which leads to the following corollary.
@@
Run Algorithm 1 $k$ times with $p$ taking each value in $[k]$ in turn, and collect all the estimates into a set $\mathcal{X}$. Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, there exists at least one estimate $\mathbf{z}\in\mathcal{X}$ that satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ when
@@
Remark 4.
While Corollary 1 guarantees the optimal sample complexity by iterating $p$ over the entire range of $[1,k]$, this approach entails running the algorithm $k$ times. To mitigate the computational cost in practice, one can adopt a logarithmic search strategy, namely, selecting candidates from a subset $\mathcal{P}=\{1,2,4,\dots,2^{\lfloor\log_{2}k\rfloor}\}$. Since the complexity bound varies smoothly with $p$, this heuristic typically identifies a parameter that yields a sample complexity within a constant factor of the global optimum, while significantly reducing the computational overhead from $O(k)$ to $O(\log k)$ rounds.
Due to computational constraints, multiple executions of the algorithm with varying $p$ are impractical. Thus, we are motivated to explore the sample complexity for specific, analytically tractable choices of $p$. A natural starting point for such an analysis is the intuitive choice of setting $p=k$. While this can be viewed as a special case of Theorem 1, we demonstrate that a more refined analysis allows us to derive a sharper result, as presented in the following corollary.
@@
Take $p=k$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number of samples $m$ satisfies
@@
Take $p=\lceil\sqrt{k}\rceil$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number $m$ of samples satisfies
@@
On one hand, the result (45) cannot be worse than (44) since
@@
Moreover, there exist some signals that make the inequality strictly hold. Therefore, (45) is indeed tighter than (44). See Section IV-B for an in-depth discussion.
On the other hand, compared to the state of the art, the results (44) and (45) are superior for certain signals. In other words, even though $s(p)$ is unknown and executing Algorithm 1 multiple times is infeasible, our algorithm taking $p=k$ or $\lceil\sqrt{k}\rceil$ can perform better compared with existing algorithms. Section IV-A provides an in-depth discussion.
@@ IV-A Broader Conditions for Optimal Sample Complexity
As discussed in Remark 1 and Remark 3, gESP offers a significant improvement over the state-of-the-art result, provided that the structural function $s(p)$ is known. This subsection further explores this superiority, demonstrating that gESP achieves the optimal sample complexity for a broader class of signals.
@@
requires $s(1)=\Theta(1)$, which is a stricter condition. The improvement results from considering $p$ entries in Step 1, thus relaxing the condition on $x_{(1)}$ to the set $\{x_{(1)},\dots,x_{(p)}\}$. Fig. 1 provides a comprehensive illustration.
@@
Moreover, even when $s(p)$ is unknown, our results (Corollary 2 and 3) can still perform better in many cases. In the following, we set $p=k$ and consider Corollary 2 for a simple discussion, although we can prove the strict superiority of the result derived in Corollary 3 (see Section IV-B). The result in Corollary 2 is
@@
Consequently, (48) outperforms the state-of-the-art result (47) once $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$. To illustrate this, we construct a parametric family of signals as follows.
@@
Consider a family of $k$-sparse signals $\mathbf{x}\in\mathbb{R}^{n}$ parameterized by a decay exponent $\delta\in(1/4,1/2)$. Let
$x_{(i)}=\begin{cases}k^{-\delta},&i=1,\\[1.99997pt] \frac{1}{2}k^{-\delta},&2\leq i\leq\lceil\sqrt{k}\rceil,\\[1.99997pt] k^{-1/2},&\lceil\sqrt{k}\rceil<i\leq k,\\[1.99997pt] 0,&i>k.\end{cases}$ (49)
For this family of signals, it holds that $s(1)=\Theta(k^{2\delta})$ whereas $s^{2}(\lceil\sqrt{k}\rceil)=\Theta(k^{4\delta-1})$, and thus we have $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$ with the improvement factor being $\Theta(k^{1-2\delta})$.
This family represents signals where most of the signal’s energy is spread fairly evenly across a long “flat tail” of many small nonzero entries, while only a small fraction sits in the first $\sqrt{k}$ larger entries, so it looks like a modest spike plus a short plateau on top of a broad, energy-dominating background. The proof for this example is provided in Appendix A-C. Interestingly, when the condition $s(\lceil{\sqrt{k}}\rceil)=\Theta(1)$ holds, our result (48) also achieves the information-theoretic bound (2). Notably, even without prior knowledge of the signal structure $s(p)$, this optimality is attainable for the same class of signals as in the case where $s(p)$ is known, simply by setting the parameter $p=k$.
@@ IV-B Strict Superiority of Corollary 3 over Corollary 2
is better than (48) in Section III-B. In this section, we provide a concrete example to illustrate this strict superiority.
@@
Assume that both $\sqrt{k}$ and $\sqrt[4]{k}$ are integers [^4] In fact, we can also construct such examples even though this assumption does not hold. However, omitting it will lead to unnecessary complications in the construction (e.g., involving terms like $\lceil\sqrt{k}\rceil$).. Construct the target signal $\mathbf{x}$ as
@@
This demonstrates the strict improvement offered by Corollary 3.
@@ V Proof
The proofs of Theorem 1 and Corollary 13 rely primarily on the analysis of each step of gESP. In the following, we establish several propositions corresponding to each step, respectively.
@@
    Provided the number of samples satisfies $m\geq Cp^{2}s^{2}(p)\log n$, it holds with probability exceeding $1-2n^{-c}$ that
@@
  • Step 2: Set $\mathbf{e}^{0}$ as the unit leading eigenvector of $\mathbf{Z}_{S^{0}}$.
@@
    Suppose the set $S^{0}$ contains $p^{\prime}$ indices and satisfies (50). Then, the inequality
    $|\mathbf{x}^{*}\mathbf{e}^{0}|\geq\frac{1}{2}\|\mathbf{x}_{S^{0}}\|$ (51)
    holds with probability at least $1-n^{-c}$ provided that $m\geq{C}{p^{\prime}s^{2}(p)}\log n$, where $c$ and $C$ are constants.
@@
    It may seem unusual to introduce a new parameter $p^{\prime}$ in this proposition, given that $S^{0}$ contains $p$ indices in Algorithm 1. However, the proposition holds even when $p^{\prime}=p$, serving as a generalization. We emphasize that this generalization will be useful in our proofs of Corollary 2 and 3, where we do not have access to the optimal choice of $p$.
  • Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in magnitude.
@@
    where $\gamma\in(0,1)$ is a constant. Moreover, suppose that (50) holds and $\mathbf{e}^{0}$ satisfies (51). Then, if the number of samples satisfies $m\geq C\gamma^{-2}{ks(p)}\log n$, it holds with probability at least $1-n^{-c}$ that
@@
    where $\mu_{1}$ is the hyperparameter given in (51) and $c_{\mu_{1},\gamma}$ is the constant depending on $\mu_{1}$ and $\gamma$. In other words, gESP selects all the indices of $S_{\gamma}$ in Step 3. Furthermore, we have
    $\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$ (52)
    Crucially, if $|x_{(k)}|$ is extremely small (violating (11)), it simply falls outside $S_{\gamma}$. This does not invalidate the proof. Instead, the “missed” small entries are absorbed into the approximation error term in Proposition 4. Since their magnitudes are small, the total error remains within the $\delta$-neighborhood ($\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$), thus removing the need for a hard lower bound on $x_{(k)}$.
  • Step 4: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
@@
    Suppose $S^{1}\supseteq S_{\gamma}$. For any constant $0<\delta<1$, Step 4 produces a signal estimate $\mathbf{z}$ falling into the $\delta$-neighborhood of $\mathbf{x}$, i.e.,
@@
    with probability exceeding $1-n^{-c}$ provided that $m\geq C\delta^{-2}k\log n$, where $c$ and $C$ are numerical constants.
We defer the proofs of Propositions 14 to the Appendix.
Now we proceed to prove the theorems in Section III.
@@
Combining the results in Proposition 14 with $p^{\prime}=p$, we can show that the sample complexity is
@@
and the overall probability is $1-n^{-c}$. Since we take the optimal value for $p$ as
@@
For each fixed $p\in[k]$, combining the results in Proposition 14 with $p^{\prime}=p$, we can show that the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-n^{-c}$ when
@@
To ensure that at least one estimate $\mathbf{z}$ satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we can apply the union bound for all $p\in[k]$. Then the sample complexity becomes
@@
and the probability is at least $1-kn^{-c}$. This is at least $1-n^{-c}$ if we take a sufficiently large constant $c$ since $n>k$. Thus the proof is complete. ∎
@@
(c) Exponential decaying signal. $k$ = 10.
Figure 3: Performance comparison of relative error and fraction of recovered support as a function of sample ratio.
@@
where $(a)$ is due to $k\geq p$, and $(b)$ holds by applying Proposition 1 to the set $S^{\prime}$ when $m$ $\geq Cp^{2}s^{2}(p)\log n$.
@@
the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-n^{-c}$. This result also holds for all $p\in[k]$ since we can take the union bound, and the probability will be $1-kn^{-c}$. It is actually $1-n^{-c}$ if we take a sufficiently large constant $c$ since $n>k$.
@@
It remains to show that (57) is equivalent to (44).
As stated following Definition 2, $ps(p)$ is monotonically increasing, and $s(p)$ is monotonically decreasing. Both functions are positive. Accordingly, we can illustrate the behavior of the two functions in (57); see Fig. 2.
@@
From Lemma 1, we know that both cases imply the same sample complexity
@@
The proof is analogous to the proof of Corollary 2. Assume any fixed $p\in[\lceil{\sqrt{k}}\rceil]$. The sample complexity for Step 1 is still $\Omega(p^{2}s^{2}(p)\log n)$, and applying Proposition 24 on the subsequent steps with $p^{\prime}=\lceil{\sqrt{k}}\rceil$ finally leads to